Jacob, Billy, and Gerald independently take a test and the probabilities that the pass the test are as follows:

Jacob: 1/5 ; Gerald:1/5 ; Billy: 1/5

Find the probability that all three pass
Answer #1

Sol)

Given

Jacob , billy and gerald independently take a test

Probability of passing the

Jacob = 1/5

Billy = 1/5

Gerald = 1/5

Probability that all three of them pass

= 1/5 * 1/5 * 1/5

= 1/125

= 0.008

= 0.8%